A metallic wire of length L is fixed between two rigid supports. If the wire is cooled through a temperature difference ∆T (Y = young's modulus, p density, a coefficients of linear expansion) then the frequency of transverse vibration is proportional to

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student Tension in the string is developed in the wire the compression due to cooling is balanced by the expansion due tension generated Thus FL/AY = Lalpha * delT F = YAalpha *delT frequency of viberation = f = 1/2L (sqrt (F/mew)f is directly proportional to (sqrt (YAalpha * delT ) )row *A*L = mew*Lf is directly proportional to sqrt (Y *alpha / row)